OpenStudy (ajanijones):

Medalss/ (30___5) Divided by 2 = 3

OpenStudy (anonymous):

So your looking for the missing numbers between 30___5?

OpenStudy (ajanijones):

Yes.

OpenStudy (anonymous):

Do you have answers to choose from?

OpenStudy (ajanijones):

No.

OpenStudy (ajanijones):

Thats why its hard.

OpenStudy (anonymous):

Can you screenshot the question?

OpenStudy (ajanijones):

Ok

OpenStudy (ajanijones):

how do i send it here

OpenStudy (ajanijones):

nvm.

OpenStudy (ajanijones):

Thats what i need help with.

OpenStudy (anonymous):

Thanks i can work with that!

OpenStudy (ajanijones):

I have already done 1.

OpenStudy (anonymous):

Your don't add numbers in that blank space you add a sign. (Add, Subtraction, Divide)

OpenStudy (ajanijones):

Ohhh....

OpenStudy (ajanijones):

Lets see.

OpenStudy (anonymous):

Hint: It's not adding or multiplying.

OpenStudy (ajanijones):

Its 3

OpenStudy (ajanijones):

I think.

OpenStudy (anonymous):

its division

OpenStudy (ajanijones):

Its division.

OpenStudy (anonymous):

Yes, you are correct.

OpenStudy (anonymous):

when u r dividing 30 by 5 u get 6...which when divided by 2 gives 3

OpenStudy (ajanijones):

There is 3 more.

OpenStudy (ajanijones):

I just have 3 4 and 5

OpenStudy (anonymous):

Ill help with three, but i think your smart enough to get through 4 and 5

OpenStudy (ajanijones):

Yea. I think i know three 1 sec.

OpenStudy (anonymous):

\[5 \frac{ 5 }{ 7 } + 2 \frac{ 1 }{ 3 }=\frac{ 169 }{ 21 }\]

OpenStudy (ajanijones):

Its Addition.

OpenStudy (anonymous):

Now change that improper fraction to a mixed number to get your answer.

OpenStudy (anonymous):

With this one they said write \[\neq or \] =

OpenStudy (anonymous):

So after you change that improper fraction to a mixed number you will see if the problem is equal or not equal